OpenStudy (anonymous):

what is 4(5z-1)-5(z+2)=2(z+1)

OpenStudy (anonymous):

Is the last term supposed to be (z = 1) or (z - 1)?

OpenStudy (anonymous):

z+1

OpenStudy (anonymous):

Oh okay. Well then.. 4(5z - 1) - 5(z + 2) = 2(z + 1) Multiply the terms using distributive property to get: 20z - 4 - (5z + 10) = 2z + 2 (the minus sign in front of the parenthesis makes them the opposite signs) 20z - 4 - 5z - 10 = 2z + 2 now move the term 2z over to the other side of the equation 20z - 4 - 5z - 10 - 2z = 2 combine like terms: (20z - 5z - 2z) + (-4 - 10) = 2 13z - 14 = 2 solve for z: 13z = 16 z = 1.2 or 16/13